OpenStudy (slinley):

_x_ + 7 = 5 4 Oh , solve for x ...

OpenStudy (lolacole12):

x=13

OpenStudy (lolacole12):

Let's solve your equation step-by-step. x+7=(5)(4) Step 1: Simplify both sides of the equation. x+7=20 Step 2: Subtract 7 from both sides. x+7−7=20−7 x=13 Answer: x=13

OpenStudy (lolacole12):

got it

OpenStudy (lolacole12):

need any more help then call me

OpenStudy (slinley):

the correct answer was -8 :} appreciate the help though , and sure.

OpenStudy (lolacole12):

Oh it was

OpenStudy (slinley):

mhm ..

OpenStudy (lolacole12):

how

OpenStudy (lolacole12):

well that's weird i did this question before and i got it write :/ Ok bye

OpenStudy (slinley):

hh

OpenStudy (slinley):

no, youre good! i mean thats what the website had... everyone solves their math problems differently .. ya know

OpenStudy (slinley):

@lolacole12